Group Selection Page HELP Controls Group Data Analysis Group ITDM WP76 DET Vacuum EEE WP78 WP74 FXE SCS SPB MID HED SQS SXP Sample Environment Operation PSPO XO TS Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2023 logbook, Contact: F.Wolff-Fabris, A.Galler, L.Samoylova, H.Sinn  Not logged in ELOG logo
Message ID: 407     Entry time: 23 Oct 2023, 05:15
Author: Raphael de Wijn 
Group: SPB 
Subject: Shift summary 

X-ray delivery

  • 13.6 keV, ~1.6 mJ and up to 352 pulses at 1.1 MHz.
  • Beam down from 10:30 to 19:00.

Optical laser delivery

  • 800 nm, 50 fs, 1 to 176 pulses at 564 kHz pumping TOPAS at 475 nm.
  • Stable delivery.

Achievements/Observations

  • PP laser and X-ray spatial and timing overlap.
  • Collected data for ~7.5h.
  • Very stable sample delivery.
ELOG V3.1.4-7c3fd00